# Gender Bias in Perception of Human Managers Extends to AI Managers

### v2

@@ 4 Discussion
Being selected for an award increased participants’ positive perceptions of the manager across all manager types. However, the magnitude of these benefits was moderated by manager gender, with male managers experiencing more pronounced gains than female managers. This aligns with pervasive societal stereotypes that link competence and leadership effectiveness with male figures [31, 19, 18]. The reinforcement of positive recognition for male managers, could reflect societal tendencies to perceive male-led successes as more legitimate or expected, a bias that some have argued is deeply ingrained in patriarchal organizational structures [44].
@@
On the other hand, negative outcomes notably harmed perceptions of female managers, particularly female-AI managers, highlighting a compounded vulnerability, where skepticism toward AI decision-making intersects with gender biases to yield harsher judgments. Female human managers also faced a notable decline in the assessment of participants, though to a lesser extent, possibly due to the perception that human managers demonstrate greater accountability and empathy [10]. By contrast, male managers exhibited greater resilience to negative outcomes, likely due to societal leniency afforded to male figures, consistent with status expectations theory [19]. This underscores the moderating role of gender in shaping perceptions of managerial decisions, illustrating a deeply embedded double standard where cultural narratives shield male leaders with presumed competence and an external locus of control, while female leaders face greater blame for unfavorable outcomes. These findings resonate with broader feminist critiques that women’s roles and societal positions are often disproportionately scrutinized within a male-dominated societal framework [45, 46].
@@
The amplified negative impact on female-AI managers emphasizes the need to address intersectional biases in workplace dynamics, where gender and technology intersect to create compounded disadvantages. Organizations aiming to integrate AI in leadership roles should consider strategies to mitigate such biases, such as transparency in AI decision-making processes and fostering a culture that values fairness and equity. Furthermore, the results point to the importance of aligning performance outcomes with how managerial decisions are perceived, as these evaluations can significantly influence trust, collaboration, and organizational morale.
Female participants and male participants showed comparable patterns in managerial perceptions, though differences also emerged. Female participants who perceive themselves as high contributors rated human managers more negatively when not awarded – possibly due to stronger expectations of fairness from human decision-makers. They also rated female managers more harshly than male participants did under non-award conditions, suggesting a more substantial in-group penalty effect, potentially driven by unmet expectations.
Male participants, when not awarded by a male human manager, rated that manager as significantly less competent and reported lower willingness to work with similar managers in the future. This may reflect that shared identity can heighten expectations, making unmet outcomes more disappointing for male participants. Additionally, postgraduate education among male participants predicted more favorable post-treatment ratings of managerial competence and fairness, suggesting that higher education may reduce negative evaluations.
@@
Our study has limitations that highlight avenues for future research. Conducted in a single setting with one specific task, our findings may not generalize to more complex scenarios involving diverse tasks, larger groups, or mixed-agent interactions between humans and AI in the same environment. Future work could expand on our design by exploring tasks across a broader spectrum of complexity, such as logic-based puzzles, creative group challenges, or multi-stage decision processes that require more sustained effort, information sharing, and interdependence among group members. This would help determine whether the observed patterns in perception and bias generalize to more complex and ecologically valid settings.
@@
Additional studies could also explore the mechanisms driving the observed biases, such as cognitive biases, stereotypes, and trust dynamics, to better understand how they shape various dimensions of perceptions, including trustworthiness, competence, fairness, and willingness to collaborate. Additionally, broader contextual factors like cultural norms, team dynamics, and prior experiences with AI or human managers could also be examined to uncover the nuanced drivers of these perceptions. These insights will be essential for designing systems across a range of settings that foster equitable, trustworthy, unbiased, and effective integration of AI in managerial and collaborative tasks.
@@
The emergence of AI in managerial and societal roles presents an opportunity to move beyond outdated frameworks and mindsets, yet it also risks perpetuating or even amplifying existing gender and sexist stereotypes if left unchecked. While significant scrutiny is placed on ensuring AI is human-centric, fair, and explainable, we must also hold ourselves accountable as humans to embody these principles in our behaviors and societal norms. History has shown how prejudice and discrimination linger in our actions and institutions, and these biases could be magnified as AI becomes more embedded in daily life. To prevent this, it is imperative to address and dismantle these stereotypes now – before they shape the interactions between humans and AI in ways that are counterproductive or harmful. Policymakers currently shaping AI regulation must consider these risks and explore bold measures, such as developing AI systems that are gender-neutral, to break free from reinforcing stereotypes. By confronting these issues today, we set the stage for a more equitable future where both humans and AI contribute to a society free from discrimination.
@@ 5 Conclusion
Our study shows that award outcomes shape perceptions toward managers, and these effects are significantly moderated by both the manager’s labeled gender and whether the manager is human or AI. Male managers benefited more from positive outcomes, while female managers, particularly female-AI, were disproportionately penalized after unfavorable outcomes. By integrating insights from social psychology, gender studies, and human-AI interaction, our work contributes to the understanding of how longstanding societal biases can manifest in emerging hybrid work environments. It highlights the nuanced ways that gender and technology intersect in shaping perceptions of managerial trust, competence, fairness, and future collaboration.
As AI systems, and particularly “AI Agents”, take on increasingly prominent roles in decision-making and leadership, addressing these compounded biases is essential. Building equitable, trustworthy, and inclusive systems will require not only fair algorithms but also a deeper understanding of how human perceptions, shaped by gender norms and societal bias, interact with algorithmic decision-makers and influence how those systems are interpreted and received.
@@ ACKNOWLEDGMENTS
Figure S10: User interface of the first round of the game.

### v3

@@ 4.4 Interaction effects: Award Outcome $\times$ Manager Type $\times$ Contribution
4.5 Participant gender differences
Female participants and male participants showed comparable patterns in managerial perceptions, though differences also emerged. Female participants who perceived themselves as high contributors rated human managers more negatively when not awarded. Male participants, when not awarded by a male human manager, rated that manager as significantly less competent and reported lower willingness to work with similar managers in the future. Additionally, postgraduate education among male participants predicted more favorable post-treatment ratings of managerial competence and fairness.
@@ 5 Discussion
The findings show that gender bias in leadership evaluations extends beyond human managers to AI decision-makers. Receiving an award increased participants’ positive perceptions of the manager across all manager types, whereas not receiving an award led to lower evaluations. However, the magnitude of these changes was moderated by manager gender. Male managers, whether human or AI, benefited more from positive outcomes and were less penalized for negative ones than female managers. The results provide new insights into how social stereotypes and expectations shape human perceptions of AI in leadership roles.
5.1 Theoretical and organizational implications
@@
Male managers exhibited greater resilience to negative outcomes, likely due to societal leniency afforded to male figures, consistent with status expectations theory [12]. The results align with pervasive societal stereotypes that link competence and leadership effectiveness with male figures [31, 12, 11]. The reinforcement of positive recognition for male managers could reflect societal tendencies to perceive male-led successes as more legitimate or expected, a bias that some have argued is deeply ingrained in patriarchal organizational structures [45].
@@
Negative outcomes notably harmed perceptions of female managers, particularly female-AI managers, highlighting a compounded vulnerability, where skepticism toward AI decision-making intersects with gender biases to yield harsher judgments. Female human managers also faced a notable decline in the assessment of participants, though to a lesser extent, possibly due to the perception that human managers demonstrate greater accountability and empathy [21]. This underscores the moderating role of gender in shaping perceptions of managerial decisions, illustrating a deeply embedded double standard where cultural narratives shield male leaders with presumed competence and an external locus of control. In contrast, female leaders face greater blame for unfavorable outcomes. These findings also resonate with broader feminist critiques that women’s roles and societal positions are often disproportionately scrutinized within a male-dominated societal framework [46, 47].
The results contribute to theory by extending role congruity and equity frameworks into the domain of AI leadership. They show that people’s perceptions of the managers are shaped by both the outcome of the decision and their gender-related expectations – even when the manager is non-human. This suggests that the same social and psychological biases that influence how we evaluate human managers also shape how we perceive artificial decision-makers.
For organizations, this has tangible implications. As AI systems take on evaluative and supervisory roles, employee trust and morale will depend on perceived fairness and social legitimacy of these systems. Training programs and workplace policies should therefore address not only algorithmic bias but also user bias – how workers interpret and respond to AI managers based on explicit or assumed gender cues.
5.2 Implications for AI leadership and design
@@
The findings show that AI managers inherit the same gendered expectations that shape human leadership judgments. The amplified negative impact on female-AI managers emphasizes the need to address intersectional biases in workplace dynamics, where gender and technology intersect to create compounded disadvantages. For designers and organizations, this underscores the importance of bias-aware AI interface design. Perceptions of AI managers depend not only on their visual or anthropomorphic cues, but also on how users interpret these features through the lens of human managerial norms and social expectations. Organizations aiming to integrate AI into leadership roles should consider strategies to mitigate such biases, such as exploring gender-neutral visual and verbal cues and fostering a culture that values fairness and equity.
5.3 Limitations
Several limitations of the current study should be acknowledged. First, the study relied on an online participant pool (Prolific) composed primarily of digitally literate U.S. adults. Such samples may not fully represent broader workforce demographics, potentially biasing perceptions toward more tech-familiar or culturally homogeneous views. Second, our findings, which were based on an experiment conducted in a single setting with a specific task, may not generalize to more complex scenarios involving diverse tasks, larger groups, or mixed-agent interactions between humans and AI in the same environment. Third, participants were aware they were in a simulated experiment, which may limit ecological validity. Field studies or hybrid human-AI team experiments in actual workplaces could enhance realism. Finally, as with many online studies, we cannot fully exclude the possibility that participants used generative AI tools (e.g., ChatGPT or Gemini). However, the nature of the task, requiring visual inspection, individual judgment, and real-time discussion, makes such use unlikely to have influenced the results.
5.4 Future directions
@@
Future work could expand on our design by exploring tasks across a broader spectrum of complexity, such as logic-based puzzles, creative group challenges, or multi-stage decision processes that require more sustained effort, information sharing, and interdependence among group members. This would help determine whether the observed patterns in perception and bias generalize to more complex and ecologically valid settings.
@@
Further research could investigate the mechanisms underlying the observed biases, such as cognitive biases, stereotypes, and trust dynamics, to gain a deeper understanding of how they influence various dimensions of perception, including trustworthiness, competence, fairness, and willingness to collaborate. Broader contextual factors, including cultural norms, team dynamics, and prior experiences with AI or human managers could also be examined to uncover the nuanced drivers of these perceptions. Future studies could employ cross-cultural samples and longitudinal designs to examine how these biases evolve and to test interventions such as anonymized AI presentation or variations in AI competence framing as potential strategies for reducing bias. Insights from such work will be essential for designing systems across a range of settings that foster equitable, trustworthy, unbiased, and effective integration of AI in managerial and collaborative tasks.
@@
The emergence of AI in managerial and societal roles presents an opportunity to move beyond outdated frameworks and mindsets, yet it also risks perpetuating or even amplifying existing gender and sexist stereotypes if left unchecked. While significant scrutiny is placed on ensuring AI is human-centric, fair, and explainable, we must also hold ourselves accountable as humans to embody these principles in our behaviors and societal norms. History has shown how prejudice and discrimination linger in our actions and institutions, and these biases could be magnified as AI becomes more embedded in daily life. To prevent this, it is imperative to address and dismantle these stereotypes now – before they shape the interactions between humans and AI in ways that are counterproductive or harmful. Policymakers currently shaping AI regulation must consider these risks and explore bold measures, such as developing AI systems that are gender-neutral, to break free from reinforcing stereotypes. By addressing these issues today, we lay the groundwork for a more equitable future where both humans and AI contribute to a society free from discrimination.
@@ 6 Conclusion
Our study reveals that award outcomes influence perceptions of managers, and these effects are significantly moderated by both the manager’s labeled gender and whether the manager is human or AI. Male managers benefited more from positive outcomes, while female managers, particularly female-AI, were disproportionately penalized after unfavorable outcomes. By integrating insights from social psychology, gender studies, and human-AI interaction, our work shows how longstanding societal biases can manifest in emerging hybrid work environments. It highlights the nuanced ways that gender and technology intersect in shaping perceptions of managerial trust, competence, fairness, and future collaboration.
Beyond documenting these patterns, the study contributes to a deeper understanding of how human social cognition transfers to AI contexts, showing that technological neutrality does not eliminate deep-seated human bias. As AI systems, and particularly “AI Agents”, take on increasingly prominent roles in decision-making and leadership, addressing these compounded biases is essential. Building equitable, trustworthy, and inclusive systems will require not only fair algorithms but also greater awareness of how human perceptions, shaped by gender norms and societal bias, interact with algorithmic decision-makers and influence how those systems are interpreted and received.
Moving forward, these insights call for a dual approach: designing AI systems that minimize gendered cues, while also fostering workplace cultures that challenge implicit bias and support equitable human-human, and human–AI collaboration. Future research should continue to explore strategies that reduce bias at both the human and system levels, ensuring that the integration of AI in leadership promotes equity rather than reproducing inequality. By addressing these challenges proactively, organizations can move toward a future where both human and AI-led decisions are evaluated fairly and constructively.
@@ ACKNOWLEDGMENTS
Figure S11: Participant view on Prolific. Clicking the link directs participants to the information sheet.
Figure S12: Information sheet and consent form (part 1).
Figure S13: Information sheet and consent form (part 2). After reading the information, participants click “I AGREE” to provide consent.
Figure S14: After providing consent, participants arrive at this page. Upon submitting their Prolific ID, they proceed to the experiment.
Figure S15: Welcome page of the experiment. Clicking “Continue” directs participants to the pre-treatment survey.
Figure S16: After completing the pre-treatment survey, participants proceed to the first round of the game.